OpenStudy (anonymous):

Write the function in vertex form, and identify its vertex. f(x) = x2 + 6x - 3 last question!

OpenStudy (amistre64):

last question on the test?

OpenStudy (anonymous):

no my last question for the day

OpenStudy (amistre64):

hmm, its hard to tell if that response is sarcastic or not lol

OpenStudy (anonymous):

its not.

OpenStudy (amistre64):

:) +9-9 to complete the sqaure then tidy it up

OpenStudy (ash2326):

We have \[f(x)=x^2+6x-3\] Let's add and subtract 9 to create square \[f(x)=x^2+6x+9-9-3\] \[f(x)=(x+3)^2-12\] Parabola's standard equation \[f(x)=a(x+h)^2+k\] the vertex is given as (-h,k) here h=3 and k=-12 so vertex is (-3, -12)

OpenStudy (amistre64):

or rather, lets add and subtract the square of half the x coeff

OpenStudy (anonymous):

thank you, ash. I see it now.

OpenStudy (ash2326):

Welcome amy

OpenStudy (amistre64):

(x+y)^2 = x^2 + 2y x + y^2 the middle term is just twice the compacted version of the square